OpenStudy (rayford):

Which of the following is the expansion of (x + 3)^5?

OpenStudy (anonymous):

alright, so first start of by making pascal's triangle. 1 (constant) 1 1 (a+b)^1 1 2 1 (a+b)^2 1 3 3 1 (a+b)^3 1 4 6 4 1 (a+b)^4 1 5 10 10 5 1 (a+b)^5

OpenStudy (rayford):

here are the answer choices: x5 + 5x4 + 10x3 + 10x2 + 5x + 1 x5 + 3x4 + 9x3 + 27x2 + 81x + 243 x5 + 15x4 + 90x3 + 270x2 + 405x + 243 None of the above

OpenStudy (anonymous):

what these mean is that is if you have \[(a+b)^n\] the expansion will be \[c_1 a^n+c_2 a^{n-1}b^{n-4}+c_3 a^{n-2}b^{n-3}+c_4 a^{n-3}b^{n-2}+c_5 a^{n-4}....... \]where the c's are the numbers on the triangle

OpenStudy (anonymous):

so fro (x+5)^5.. you have x=a, and b=5 lookingat the triangle for the 5th power

OpenStudy (rayford):

so it's none of the above obviously?

OpenStudy (rayford):

am i correct?

OpenStudy (anonymous):

so your first is going to x^5

OpenStudy (anonymous):

second term is \[5x^4*3=5x^4(b)=15x^4\]

OpenStudy (anonymous):

next term is \[10x^{n-2}b{n-3}=10x^3*3^2=90x^3\]